# Physics-Inspired Extrapolation for efficient error mitigation and hardware certification

### v2

@@ III Results
The PIE method enables accurate and efficient error mitigation without requiring explicit noise characterization. In this section, we focus on benchmarking PIE against the standard ZNE techniques and report the corresponding results. As a proof-of-concept, we primarily use the transverse-field Ising model to benchmark our results. The main reason behind using this model is that it is one of the simplest yet non-trivial many-body models exhibiting quantum phase transition [52, 53], is exactly solvable in one-dimension [54, 55], and captures the physics of some real magnetic materials [56, 57, 58, 59, 60, 61, 62, 63]. We also benchmark PIE using quantum chemistry simulations, in particular, by calculating the ground-state energies of the hydrogen (H2) and lithium hydride (LiH) molecules. These results (and corresponding experimental setup) are presented in the subsections below. The application of PIE in much more complex chemical molecular systems will be shown in our subsequent study [64].
@@ III.1 Ising model with transverse-field
with $N$ representing the total number of spins or qubits, and $\Gamma=0.5$ being the transverse field strength. The $\sigma$’s are the usual Pauli matrices. The time-evolution is approximated via the first-order Suzuki-Trotter decomposition,
@@
$M_{z}(t)=\sum_{i=0}^{N}\dfrac{{\rm Tr}[\sigma^{z}_{i}U_{R}|\Psi(0)\rangle\langle\Psi(0)|U_{R}^{\dagger}]}{N}\,$, (12)
given the noisy hardware and various QEM techniques at hand. We compute its evolution at different times $t$, starting from an initial state in which all spins are aligned in the positive $\sigma_{z}$ direction such that $M_{z}(0)=1$. While the one-dimensional Ising model described in Eq. (9) admits analytical solutions for an arbitrary number of spins [65, 66], our primary focus is not on the simulation error introduced by the Trotter approximation in Eq. (10). Instead, we focus on evaluating the performance of error mitigation algorithms when applied to actual hardware demonstrations and theoretical studies under realistic and commonly encountered noise models, such as depolarizing, dephasing, and random Pauli noises. In this respect, the noise-free output of the Trotter circuit in Eq. (10) is taken as the ideal result regardless of the number of Trotter steps. We benchmark the PIE method against different extrapolation-based error mitigation techniques by deploying them to estimate the global transverse magnetization for different evolution times. The noise amplification is carried out by circuit folding [17], where the trotterized time-evolution operator Eq. (11) is replaced by
@@
Figure 5: (Color online) Numerical results of the global magnetization $M_{z}$ on the three Trotter steps simulation of the one-dimensional Ising chain at time $t=1.5$ with N = 8 spins using a depolarizing noise model. We display the error mitigated results (dots) from PIE, exponential ZNE, and polynomial quadratic ZNE techniques, together with the standard deviation of the results (bars). We also show the unmitigated or noisy result and the noise-free simulation, with the corresponding standard deviation (bars and shaded area, respectively). All experiments were performed using four extrapolation points and 4096 shots.
@@ III.2 Quantum Chemistry example
where $h_{pq}$ are one-particle integrals given by
$h_{pq}=\int d\sigma\phi^{*}_{p}(\sigma)\left(\frac{\nabla^{2}_{r}}{2}-\sum_{i}\frac{Z_{i}}{|R_{i}-r|}\right)\phi_{q}(\sigma)\,$, (26)
@@
and $a^{\dagger}_{i}(a_{i})$ is the creation (annihilation) operator acting on $i$-th state.
Figure 6: (Color online) Experimental error-mitigated results on the ground-state energy calculation of two distinct molecules. The experiments were performed on the IBM Eagle superconducting processor ibm sherbrooke, using 4096 shots of a hardware efficient SU(2) 2-local ansatz with two layer repetitions and optimal angles. Together with the mitigated result, we show the PIE extrapolation from the noise amplified expectation values, the result from a noise-free execution of the quantum circuit, and the topology of the quantum computer with the qubits selected for the experiments in dark blue. a) Hydrogen molecule H2 with an equilibrium distance of 0.74 Å (four qubit ansatz). b) Lithium hydride LiH with an equilibrium distance of 1.6 Å (twelve qubit ansatz).
We chose the STO-3G basis set and the equilibrium distance between two hydrogen atoms of 0.74 Å and distance between Li and H as 1.60 Å, and performed Hartree-Fock calculations to get $h_{pq}$ and $h_{pqrs}$ using PySCF [67, 68]. We applied the Jordan-Wigner transformation to map the second-quantized electronic Hamiltonian in (25) to qubit Hamiltonian ($\hat{H}_{q}$) which can be written in the form below
@@
We used the variational quantum eigensolver on a local simulator (a classical computer) to obtain the optimal parameters ${\theta}_{opt}$} of a parameterized quantum circuit, $|\Psi(\theta)\rangle=|U(\theta)\rangle$, that minimize the energy of the molecules ($E=\langle\Psi(\theta)|\hat{H}_{q}|\Psi(\theta)\rangle$). We employed a hardware-efficient SU(2) 2-local ansatz with two layers (each layer consisting of Pauli Y and Pauli Z parameterized single-qubit rotations and two-qubit CNOT entangling gates between nearest neighbors), and performed parameter update using the classical optimizer COBYLA [69]. We then calculate the molecule energy using 4096 shots of the quantum circuit with optimal parameters, $|\Psi\rangle\equiv|\Psi(\theta_{opt})\rangle$, executed on the IBM Eagle superconducting processor, ibm sherbrooke. The 4- and 12-qubit quantum circuits for the H2 and LiH molecules lead to a depth of 25 and 53, respectively, after transpilation. Note that we are not interested in the capability of the hardware-efficient ansatz to approximate the ground state of the molecule, but in testing how PIE is able to compute the expected value of a noise-free execution of the ansatz with the most optimal parameters.
For the quantum circuit and the Hamiltonian that we prepared, we calculated the expectation value, $E=\langle\Psi|\hat{H}_{q}|\Psi\rangle$, and its variance, $\text{Var}=\langle\Psi|\hat{H}_{q}^{2}|\Psi\rangle-\langle\Psi|\hat{H}_{q}|\Psi\rangle^{2}$, at different noise levels ($\lambda=1,3,5$) to perform PIE. We found that PIE provides good error-mitigated results as shown in Fig. 6. In particular, in the case of the 4 qubit H2 molecule, the reference value electronic ground state energy is -1.852 Ha, the noise-free execution of the hardware-efficient ansatz with optimal parameters result is $1.855\pm 0.005$ Ha, the experimental unmitigated result ($\lambda=1$) is $1.793\pm 0.004$ Ha, and the PIE mitigated result is $1.86\pm 0.01$ Ha. For the 12 qubit LiH molecule, the reference value is -8.875 Ha, the noise-free result is $8.85\pm 0.04$ Ha, the experimental unmitigated result is $7.90\pm 0.02$ Ha, and the PIE mitigated result is $8.8\pm 0.4$ Ha.
@@ IV Discussion
In this paper, we propose a linear runtime extrapolation-based error mitigation protocol built on the EMRE framework. Unlike purely phenomenological data fitting approaches, PIE’s extrapolation fit is operationally guided from EMRE, without requiring explicit hardware noise characterization. This operational grounding also explains why the parameters of heuristic ZNE approaches cannot be assigned physically meaningful interpretations or used for tasks such as circuit certification.
@@
Through both numerical simulations and experimental demonstrations, we show that the protocol performs optimally in low-noise regimes where early FTQC algorithms are expected to operate or for short-depth circuits. Moreover, PIE provides a clear physical interpretation to the experimental data fitting process and its associated parameters, a feature that is often absent in extrapolation-based mitigation techniques. Specifically, the slope of the fit approximates the max-relative entropy, a measure that quantifies the distance between the ideal and the noisy quantum circuit. This information enables users to assess the impact of noise on their circuits and determine whether the collected data remains meaningful. In doing so, PIE not only extracts physically relevant quantities that characterize error mechanisms, but also facilitates a form of hardware certification where the slope serves as a proxy for measuring the deviation between implemented and ideal circuit behavior. This dual capability fundamentally distinguishes PIE from alternative mitigation methods such as ZNE, which rely solely on empirical fitting without offering insight into the underlying error structure or hardware quality. Since PIE’s extrapolation function is derived from the first principles under physical assumptions involving max-relative entropy, this level of interpretability cannot, in general, be obtained from heuristically motivated ZNE fits.
@@
From an operational perspective, the PIE method offers efficient error mitigation with enhanced robustness and notably reduced variance compared to non-linear exponential fitting techniques. This enhanced stability and variance stems from the linear extrapolation model enabled by the restricted evolution framework of EMRE.
@@
Looking ahead, a key direction is to extend this protocol to handle higher-noise settings and integrating it with fault-tolerance schemes to provide a scalable pathway toward reliable quantum computation. The challenge will be to design such methods to remain fully operationally meaningful rather than heuristic. Moreover, beyond full circuit folding, PIE may benefit from the use of more accurate noise amplification techniques, such as partial folding, analog noise strength control [14], or circuit quantum unoptimization [70]. It is also interesting to see how PIE can be used as an efficient computational tool to approximate max relative entropy in resource theoretic tasks. Another open avenue of research is to develop efficient methods to compute the exact distance between the ideal and erroneous circuit, further strengthening the interpretability and diagnostic power of error mitigation.
@@ Acknowledgments
S5 Inverse EMRE technique
@@
S5.1 Method
Error mitigation by Restricted Evolution (EMRE) works by restricting the evolution of the input quantum state [26]. The output obtained from this restricted evolution gives the estimate of the ideal expectation value. EMRE trades off sampling efficiency with a small non-zero bias. The restriction is performed in such a way that sampling the circuit is efficient. In Ref. [26], two of the authors of this paper introduced the concept of a generalized quasi-probabilistic decomposition (GQPD) for a unitary channel $\mathcal{U}$. This decomposition is given by
$\mathcal{U}=s\mathcal{B}-(s-1)\mathcal{N}\,$, (S1)
@@
The above protocol relies on obtaining a generalized QPD for each gate in the circuit. The smallest bias is achieved when the generalized QPD used is optimal. In [26], it has been shown that finding the optimal generalized QPD can be cast as an optimization problem. This optimization problem requires a characterization of the noise in the hardware to characterize the set of noisy operations, making it challenging to obtain the optimal solution. Thus, with unknown hardware noise, it seems difficult to practically implement and mitigate noise using EMRE.
@@
$G\approx s\left(\sum_{i}p_{i}P_{i}\right)\circ G^{\rm err}\,.$ (S3)
@@
$R^{\Lambda}=\dfrac{1}{2^{N}}{\rm Tr}\left[P_{i}\Lambda(P_{j})\right]$ (S4)
with $\Lambda$ being the one- ($N=1$) or two-qubit ($N=2$) quantum channels ${\mathcal{E}}^{-1}$, $G$ or $G^{\rm err}$, and $P_{i}$ are the elements is the $N$-qubit Pauli basis ($\{I,X,Y,Z\}$ for $N=1$ and $\{II,IX,IY,IZ,XI,XX,XY,XZ,YI,YX,YY,YZ,ZI,ZX,ZY,ZZ\}$ for $N=2$). In Fig. S7 and Fig. S8 we show some examples of erroneous gate PTMs, $R^{G^{\rm err}}$, and ideal gate PTMs, $R^{G}$, obtained from numerical simulations and Eq. (S4). Assuming random Pauli noise, the PTM representation of the inverse noise channel ${R^{\mathcal{E}}}^{-1}$ together with the PTM representation of erroneous Pauli gates (where $\Lambda=P_{i}$) allows for the calculation of a GQPD $\{q^{G}_{i}\}$ by solving an optimization problem as given in Ref. [26], so that ${R^{\mathcal{E}}}^{-1}\approx\sum_{i}q^{G}_{i}P_{i}$ and $\sum_{i}q^{G}_{i}=1$. Recall that a QPD is also a GQPD. In Fig. S9 we show examples of QPDs of the inverse noise operations associated with the single-qubit X gate and the two-qubit controlled-NOT gate in numerical simulations.
Now, to implement the inverse EMRE technique, we restrict the QPD to its positive part. The normalized coefficients of the positive part represent the probabilities $\{p^{G}_{i}=q^{G}_{i}/s^{G}|q^{G}_{i}>0\}_{i}$ of the occurrence of Pauli gates $\{P_{i}\}_{i}$ such that each ideal quantum gate can be approximated by sampling from the $N$-qubit Pauli basis $\{P_{i}\}$ with probabilities $p^{G}_{i}$ as depicted in Eq. (S3). The robustness $s^{G}$ of the gate $G$ is the sum of positive coefficients of the QPD, i.e., $s=\sum_{i}q^{G}_{i}$. In practice, a number of circuits $K$ is sampled where the gates are probabilistically chosen following Eq. (S3), the expectation value of the observable of interest is computed for each sampled circuit $\langle\mathcal{O}\rangle_{k}$ with $k\in[1,2,...,K]$, and the mitigated estimate is calculated as the average of those expectation values,
$\langle\mathcal{O}\rangle_{\textnormal{mitigated}}=\frac{s}{K}\sum_{k=1}^{K}\langle\mathcal{O}\rangle_{k}\;\;,$ (S5)
where $s=\prod_{G}s^{G}$. The inverse technique to perform Probabilistic Error Cancellation (PEC) follows similarly with $\{p^{(PEC)G}_{i}\}=\{|q^{G}_{i}|\}$ and $s^{(PEC)G}=\sum_{i}p^{(PEC)G}_{i}$. Contrary to EMRE, PEC provides an unbiased result provided a sufficiently large number of circuit samples, however, it has been shown that EMRE achieves more accurate estimates in a scenario with reduced number of samples. In [26], the Hybrid Error Mitigation By Restricted Evolution (HEMRE) is also introduced, where only a subset of gates is restricted to positive QPDs. In particular, we consider an HEMRE protocol where $\{p^{(HEMRE)G}_{i}\}=\{q^{G}_{i}|q^{G}_{i}>0\}$ if $G$ is a single-qubit gate, $\{p^{(HEMRE)G}_{i}\}=\{|q^{G}_{i}|\}$ if $G$ is a two-qubit gate, and $s^{(HEMRE)G}=\sum_{i}p^{(HEMRE)G}_{i}$.
@@
$\hat{H}=-\sum_{i=0}^{N-2}\sigma^{x}_{i}\sigma^{x}_{i+1}-\sigma^{x}_{N-1}\sigma^{x}_{0}-\Gamma\sum_{i=0}^{N-1}\sigma^{z}_{i}\,$, (S6)
with $N$ the number of qubits, and $\Gamma=0.5$ the transverse field strength. The dynamics is approximated via the first-order Trotter-Suzuki decomposition using one Trotter step to compute the global transverse magnetization at time $t=0.5$. We test the three protocols using the described inverse technique with a constant number of sampled circuits $K=1200$. In Figure S10 and S11 we show the numerical and experimental results respectively. Note that the time between the QPD calculation and the sampling must be as small as possible to reduce errors coming from physical fluctuations occurring in the quantum computer.
@@
Figure S7: (Color online) Examples of the Pauli transfer matrices representation (PTMs) of the identity (I), Pauli-X (X), Sqrt-X (SX), and rotation of angle $\pi$ around the Z axis (RZ) single-qubit gates. (a) Erroneous gate PTMs obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots, (b) Ideal gate PTMs obtained from noise-free statevector simulator.
@@
Figure S8: (Color online) Examples of the Pauli transfer matrices representation (PTMs) of the CNOT gate. (left) Erroneous gate PTM obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots, (right) Ideal gate PTM obtained from noise-free statevector simulator.
@@
Figure S9: (Color online) Quasi-probabilistic decompositions (QPDs) of the inverse noise operations assuming random Pauli noise of (a) the experimental single-qubit gate X, (b) the experimental two-qubit gate CNOT. The QPDS were obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots.

### v3

@@ III Results
The PIE method enables accurate and efficient error mitigation without requiring explicit noise characterization. In this section, we focus on benchmarking PIE against the standard ZNE techniques and report the corresponding results. As a proof-of-concept, we primarily use the transverse-field Ising model to benchmark our results. The main reason behind using this model is that it is one of the simplest yet non-trivial many-body models exhibiting quantum phase transition Sachdev (2011); Dutta et al. (2015), is exactly solvable in one-dimension Pfeuty (1970); Heyl (2014), and captures the physics of some real magnetic materials Coldea et al. (2010); Chen (2019); Cui et al. (2019); Wang et al. (2018); Matsuda et al. (2017); Faure et al. (2018); Okutani et al. (2015); Kyaw et al. (2020). We also benchmark PIE using quantum chemistry simulations, in particular, by calculating the ground-state energies of the hydrogen (H2) and lithium hydride (LiH) molecules. These results (and corresponding experimental setup) are presented in the subsections below. The application of PIE in much more complex chemical molecular systems will be shown in our subsequent study Cabral et al. (2025).
@@ III.1 Ising model with transverse-field
with $N$ representing the total number of spins or qubits, $\Gamma=0.5$ being the transverse field strength, and $\sigma^{x}$, $\sigma^{z}$ denoting the usual Pauli matrices. The time-evolution is approximated via the first-order Suzuki-Trotter decomposition,
@@
$M_{z}(t)=\sum_{i=0}^{N-1}\dfrac{{\rm Tr}[\sigma^{z}_{i}U_{R}|\Psi(0)\rangle\langle\Psi(0)|U_{R}^{\dagger}]}{N}\,$, (12)
given the noisy hardware and various QEM techniques at hand. We compute its evolution at different times $t$, starting from an initial state in which all spins are aligned in the positive $\sigma_{z}$ direction such that $M_{z}(0)=1$. While the one-dimensional Ising model described in Eq. (9) admits analytical solutions for an arbitrary number of spins Suzuki et al. (2013); Auerbach (1994), our primary focus is not on the simulation error introduced by the Trotter approximation in Eq. (10). Instead, we focus on evaluating the performance of error mitigation algorithms when applied to actual hardware demonstrations and theoretical studies under realistic and commonly encountered noise models, such as depolarizing, dephasing, and random Pauli noises. In this respect, the noise-free output of the Trotter circuit in Eq. (10) is taken as the ideal result regardless of the number of Trotter steps. We benchmark the PIE method against different extrapolation-based error mitigation techniques by deploying them to estimate the global transverse magnetization for different evolution times. The noise amplification is carried out by circuit folding Giurgica-Tiron et al. (2020), where the trotterized time-evolution operator Eq. (11) is replaced by
@@
Figure 5: (Color online) Numerical results of the global magnetization $M_{z}$ on the three Trotter steps simulation of the one-dimensional Ising chain at time $t=1.5$ with N = 8 spins using a depolarizing noise model. We display the error mitigated results (dots) from three extrapolation techniques (PIE, exponential, and polynomial quadratic), together with the standard deviation of the results (bars). We also show the unmitigated or noisy result and the noise-free simulation, with the corresponding standard deviation (bars and shaded area, respectively). All experiments were performed using four extrapolation points and 4096 shots.
@@ III.2 Quantum Chemistry example
where $a^{\dagger}_{i}(a_{i})$ is the creation (annihilation) operator acting on $i$-th state, $h_{pq}$ are one-particle integrals given by
$h_{pq}=\int d\sigma\phi^{*}_{p}(\sigma)\left(-\frac{\nabla^{2}_{r}}{2}-\sum_{i}\frac{Z_{i}}{|R_{i}-r|}\right)\phi_{q}(\sigma)\,$, (26)
@@
$\phi_{i}$ is a basis used to represent the wave function, and $\sigma_{i}$ is a spatial and spin coordinate with $\sigma_{i}=(r_{i};s_{i})$ Cao et al. (2019). The first and second terms in parentheses in Eq. (26) represent the kinetic and potential energy of electrons, respectively.
Figure 6: (Color online) Experimental error-mitigated results on the ground-state energy calculation of two distinct molecules. The experiments were performed on the IBM Eagle superconducting processor ibm sherbrooke, using 4096 shots of a hardware-efficient SU(2) 2-local ansatz with two-layer repetitions and optimal angles. Together with the mitigated result, we show the PIE extrapolation from the noise amplified expectation values, the result from a noise-free execution of the quantum circuit, and the topology of the quantum computer with the qubits selected for the experiments in dark blue. a) Hydrogen molecule H2 with an equilibrium distance of 0.74 Å (four qubit ansatz). b) Lithium hydride LiH with an equilibrium distance of 1.6 Å (twelve qubit ansatz).
We chose the STO-3G basis set and the equilibrium distance between two hydrogen atoms of 0.74 Å and distance between Li and H as 1.60 Å, and performed Hartree-Fock calculations to get $h_{pq}$ and $h_{pqrs}$ using PySCF Sun et al. (2018, 2020). We applied the Jordan-Wigner transformation to map the second-quantized electronic Hamiltonian in (25) to a qubit Hamiltonian ($\hat{H}_{q}$), which can be written in the form below
@@
We used the variational quantum eigensolver on a local simulator (a classical computer) to obtain the optimal parameters ${\theta}_{opt}$} of a parameterized quantum circuit, $|\Psi(\theta)\rangle=|U(\theta)\rangle$, that minimize the energy of the molecules ($E=\langle\Psi(\theta)|\hat{H}_{q}|\Psi(\theta)\rangle$). We employed a hardware-efficient SU(2) 2-local ansatz with two layers (each layer consisting of Pauli Y and Pauli Z parameterized single-qubit rotations and two-qubit CNOT entangling gates between nearest neighbors), and performed parameter update using the classical optimizer COBYLA Powell (1994). We then calculate the molecule energy using 4096 shots of the quantum circuit with optimal parameters, $|\Psi\rangle\equiv|\Psi(\theta_{opt})\rangle$, executed on the IBM Eagle superconducting processor, ibm sherbrooke. The 4- and 12-qubit quantum circuits for the H2 and LiH molecules lead to a depth of 25 and 53, respectively, after transpilation. Note that we are not interested in the capability of the hardware-efficient ansatz to approximate the ground state of the molecule, but in testing how PIE is able to compute the expected value of a noise-free execution of the ansatz with the most optimal parameters. Moreover, unlike the global transverse-field magnetization in Eq. (12), the energy of a fully mixed state for molecular Hamiltonians is generally non‑zero. This arises from the nuclear–nuclear repulsion term as well as the non‑vanishing trace of the electronic Hamiltonian. Consequently, since Eq. (16) asymptotically approaches zero, the PIE fit should be regarded as an approximation in the short $\lambda$ regime. In general, to work with a non-traceless Hamiltonian, one must shift the qubit Hamiltonian (as in Eq. (28)) to the traceless part, obtain results using PIE and then shift it back to obtain the estimate of the ideal result.
For the quantum circuit and the Hamiltonian that we prepared, we calculated the expectation value, $E=\langle\Psi|\hat{H}_{q}|\Psi\rangle$, and its variance, $\text{Var}=\langle\Psi|\hat{H}_{q}^{2}|\Psi\rangle-\langle\Psi|\hat{H}_{q}|\Psi\rangle^{2}$, at different noise levels ($\lambda=1,3,5$) to perform PIE. We found that PIE provides good error-mitigated results as shown in Fig. 6. In particular, in the case of the 4 qubit H2 molecule, the reference value electronic ground state energy is -1.852 Ha, the noise-free execution of the hardware-efficient ansatz with optimal parameters result is $1.855\pm 0.005$ Ha, the experimental unmitigated result ($\lambda=1$) is $1.793\pm 0.004$ Ha, and the PIE mitigated result is $1.86\pm 0.01$ Ha. For the 12-qubit LiH molecule, the reference value is -8.875 Ha, the noise-free result is $8.85\pm 0.04$ Ha, the experimental unmitigated result is $7.90\pm 0.02$ Ha, and the PIE mitigated result is $8.8\pm 0.4$ Ha.
@@ IV Discussion
In this paper, we propose a linear runtime extrapolation-based error mitigation protocol built on the EMRE framework. Unlike purely phenomenological data-fitting approaches, PIE’s extrapolation fit is operationally guided from EMRE, without requiring explicit hardware noise characterization. This derivation singles out a specific point in the full ZNE parameter space and also provides a clear physical meaning to the parameters in the fit. In this sense, PIE is not a simplified exponential ZNE, but a physically inspired specialization of ZNE where the extrapolation fit is derived rather than selected heuristically. Furthermore, this operational grounding also explains why the parameters of heuristic ZNE approaches cannot be assigned physically meaningful interpretations or used for tasks such as circuit certification.
@@
Through both numerical simulations and experimental demonstrations, we show that the protocol performs optimally in low-noise regimes where early FTQC algorithms are expected to operate or for short-depth circuits. Moreover, PIE provides a clear physical interpretation to the experimental data fitting process and its associated parameters, a feature that is often absent in extrapolation-based mitigation techniques. Specifically, the slope of the fit approximates the max-relative entropy, a measure that quantifies the distance between the ideal and the noisy quantum circuit. This information enables users to assess the impact of noise on their circuits and determine whether the collected data remains meaningful. In doing so, PIE not only extracts physically relevant quantities that characterize error mechanisms, but also facilitates a form of hardware certification where the slope serves as a proxy for measuring the deviation between implemented and ideal circuit behavior. This dual capability distinguishes PIE from alternative mitigation methods such as conventional ZNE, which rely solely on empirical fitting without offering insight into the underlying error structure or hardware quality. Since PIE’s extrapolation function is derived from the first principles under physical assumptions involving max-relative entropy, this level of interpretability cannot, in general, be obtained from heuristically motivated ZNE fits.
@@
From an operational perspective, the PIE method offers efficient error mitigation with enhanced robustness and notably reduced variance compared to non-linear exponential fitting techniques. This enhanced stability and variance stems from the linear extrapolation model enabled by the restricted evolution framework of EMRE. Furthermore, unlike recent learning-based methods, PIE does not require a pre-trained model or a large amount of data, while offering robust performance in the low-noise regime relevant to PIE compared to methods such as Clifford Data Regression (see Supp. Info. Sec. 4).
@@
Looking ahead, a key direction is to extend this protocol to handle higher-noise settings and integrating it with fault-tolerance schemes to provide a scalable pathway toward reliable quantum computation. The challenge will be to design such methods to remain fully operationally meaningful rather than heuristic. Another direction for generalization is to consider the case of variable noise throughout the circuit. To tackle variable noise, one approach can be to fold gates at their occurrence. In this way, one can have a higher confidence that the inverse unitary gate is not experiencing a different noise than the original unitary gate itself. Adaptive folding Koenig et al. (2025), use of purity Jin et al. (2024) and other approaches Koenig et al. (2024); Hour et al. (2024) were recently proposed for ZNE to tackle this practical problem; the question would be to incorporate it with a non-heuristically motivated extrapolation fit such as PIE. Moreover, beyond full circuit folding, PIE may benefit from the use of more accurate noise amplification techniques, such as partial folding, analog noise strength control Temme et al. (2017), or circuit quantum unoptimization Pelofske and Russo (2025). The combination of PIE with randomized compiling can also improve the performance of the method in the presence of coherent errors Kurita et al. (2022). It is also interesting to see how PIE can be used as an efficient computational tool to approximate max relative entropy in resource theoretic tasks. Another open avenue of research is to develop efficient methods to compute the exact distance between the ideal and erroneous circuit, further strengthening the interpretability and diagnostic power of error mitigation.
@@ Acknowledgments
Appendix A Inverse EMRE technique
@@
A.1 Method
Error mitigation by Restricted Evolution (EMRE) works by restricting the evolution of the input quantum state Saxena and Kyaw (2024). The output obtained from this restricted evolution gives the estimate of the ideal expectation value. EMRE trades off sampling efficiency with a small non-zero bias. The restriction is performed in such a way that sampling the circuit is efficient. In Ref. Saxena and Kyaw (2024), two of the authors of this paper introduced the concept of a generalized quasi-probabilistic decomposition (GQPD) for a unitary channel $\mathcal{U}$. This decomposition is given by
$\mathcal{U}=s\mathcal{B}-(s-1)\mathcal{N}\,$, (1)
@@ A.1 Method
The above protocol relies on obtaining a generalized QPD for each gate in the circuit. The smallest bias is achieved when the generalized QPD used is optimal. In Saxena and Kyaw (2024), it has been shown that finding the optimal generalized QPD can be cast as an optimization problem. This optimization problem requires a characterization of the noise in the hardware to characterize the set of noisy operations, making it challenging to obtain the optimal solution. Thus, with unknown hardware noise, it seems difficult to practically implement and mitigate noise using EMRE.
@@
$G\approx s\left(\sum_{i}p_{i}P_{i}\right)\circ G^{\rm err}\,.$ (3)
@@ A.2 Numerical analysis
$R^{\Lambda}=\dfrac{1}{2^{N}}{\rm Tr}\left[P_{i}\Lambda(P_{j})\right]$ (4)
with $\Lambda$ being the one- ($N=1$) or two-qubit ($N=2$) quantum channels ${\mathcal{E}}^{-1}$, $G$ or $G^{\rm err}$, and $P_{i}$ are the elements is the $N$-qubit Pauli basis ($\{I,X,Y,Z\}$ for $N=1$ and $\{II,IX,IY,IZ,XI,XX,XY,XZ,YI,YX,YY,YZ,ZI,ZX,ZY,ZZ\}$ for $N=2$). In Fig. 7 and Fig. 8 we show some examples of erroneous gate PTMs, $R^{G^{\rm err}}$, and ideal gate PTMs, $R^{G}$, obtained from numerical simulations and Eq. (4). Assuming random Pauli noise, the PTM representation of the inverse noise channel ${R^{\mathcal{E}}}^{-1}$ together with the PTM representation of erroneous Pauli gates (where $\Lambda=P_{i}$) allows for the calculation of a GQPD $\{q^{G}_{i}\}$ by solving an optimization problem as given in Ref. Saxena and Kyaw (2024), so that ${R^{\mathcal{E}}}^{-1}\approx\sum_{i}q^{G}_{i}P_{i}$ and $\sum_{i}q^{G}_{i}=1$. Recall that a QPD is also a GQPD. In Fig. 9 we show examples of QPDs of the inverse noise operations associated with the single-qubit X gate and the two-qubit controlled-NOT gate in numerical simulations.
Now, to implement the inverse EMRE technique, we restrict the QPD to its positive part. The normalized coefficients of the positive part represent the probabilities $\{p^{G}_{i}=q^{G}_{i}/s^{G}|q^{G}_{i}>0\}_{i}$ of the occurrence of Pauli gates $\{P_{i}\}_{i}$ such that each ideal quantum gate can be approximated by sampling from the $N$-qubit Pauli basis $\{P_{i}\}$ with probabilities $p^{G}_{i}$ as depicted in Eq. (3). The robustness $s^{G}$ of the gate $G$ is the sum of positive coefficients of the QPD, i.e., $s=\sum_{i}q^{G}_{i}$. In practice, a number of circuits $K$ is sampled where the gates are probabilistically chosen following Eq. (3), the expectation value of the observable of interest is computed for each sampled circuit $\langle\mathcal{O}\rangle_{k}$ with $k\in[1,2,...,K]$, and the mitigated estimate is calculated as the average of those expectation values,
$\langle\mathcal{O}\rangle_{\textnormal{mitigated}}=\frac{s}{K}\sum_{k=1}^{K}\langle\mathcal{O}\rangle_{k}\;\;,$ (5)
where $s=\prod_{G}s^{G}$. The inverse technique to perform Probabilistic Error Cancellation (PEC) follows similarly with $\{p^{(PEC)G}_{i}\}=\{|q^{G}_{i}|\}$ and $s^{(PEC)G}=\sum_{i}p^{(PEC)G}_{i}$. Contrary to EMRE, PEC provides an unbiased result provided a sufficiently large number of circuit samples, however, it has been shown that EMRE achieves more accurate estimates in a scenario with reduced number of samples. In Saxena and Kyaw (2024), the Hybrid Error Mitigation By Restricted Evolution (HEMRE) is also introduced, where only a subset of gates is restricted to positive QPDs. In particular, we consider an HEMRE protocol where $\{p^{(HEMRE)G}_{i}\}=\{q^{G}_{i}|q^{G}_{i}>0\}$ if $G$ is a single-qubit gate, $\{p^{(HEMRE)G}_{i}\}=\{|q^{G}_{i}|\}$ if $G$ is a two-qubit gate, and $s^{(HEMRE)G}=\sum_{i}p^{(HEMRE)G}_{i}$.
@@
$\hat{H}=-\sum_{i=0}^{N-2}\sigma^{x}_{i}\sigma^{x}_{i+1}-\sigma^{x}_{N-1}\sigma^{x}_{0}-\Gamma\sum_{i=0}^{N-1}\sigma^{z}_{i}\,$, (6)
with $N$ the number of qubits, and $\Gamma=0.5$ the transverse field strength. The dynamics is approximated via the first-order Trotter-Suzuki decomposition using one Trotter step to compute the global transverse magnetization at time $t=0.5$. We test the three protocols using the described inverse technique with a constant number of sampled circuits $K=1200$. In Figure 10 and 11 we show the numerical and experimental results respectively. Note that the time between the QPD calculation and the sampling must be as small as possible to reduce errors coming from physical fluctuations occurring in the quantum computer.
@@
Figure 7: (Color online) Examples of the Pauli transfer matrices representation (PTMs) of the identity (I), Pauli-X (X), Sqrt-X (SX), and rotation of angle $\pi$ around the Z axis (RZ) single-qubit gates. (a) Erroneous gate PTMs obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots, (b) Ideal gate PTMs obtained from noise-free statevector simulator.
@@
Figure 8: (Color online) Examples of the Pauli transfer matrices representation (PTMs) of the CNOT gate. (left) Erroneous gate PTM obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots, (right) Ideal gate PTM obtained from noise-free statevector simulator.
@@
Figure 9: (Color online) Quasi-probabilistic decompositions (QPDs) of the inverse noise operations assuming random Pauli noise of (a) the experimental single-qubit gate X, (b) the experimental two-qubit gate CNOT. The QPDS were obtained from numerical simulations based on a noise model provided by qiskit that mimics the estimated noise of the IBM Eagle superconducting processor ibm kyiv using 4096 shots.